The amount of leaf protein (mg/g fresh weight) in soybean plants of a particular variety was determined for a sample of six plants, yielding the following data: 12.9, 15.5, 13.2, 7.5, 6.1, 8.9 Compute: The standard deviation
Answer #1

Answer:

The given data is   12.9, 15.5, 13.2, 7.5, 6.1, 8.9

x (x-x̅ ) (x - x̅)²
12.9 2.21667 4.913626
15.5 4.81667 23.20031
13.2 2.51667 6.333628
7.5 -3.18333 10.13359
6.1 -4.58333 21.00691
8.9 -1.78333 3.180266
Total = 64.1 68.76833

Mean = ∑x/n         =64.1/6     =10.68333

Sample standard deviation = √(∑(x - x̅)²/n-1)     =√(68.76833/6-1)     =3.708593534     =3.7086

Sample Standard Deviation = 3.7086